• Main INDEX
  • Monthly INDEX
  • PREV
  • NEXT

    User name Z. Chai

    Log entry time 00:10:06 on June27,2000

    Entry number 44942

    keyword=Swing shift summary

    16:00 FPP chamber still has problems. Ask for access to check it
    There was the gas flow problem. The E-Arm VDC efficiency was only 90%. Both the E and H-Arm VDC multihit distributrion were not good. Checked
    the FPP TDC output, seemed no problem. Jack fixed the gas problem.
    18:00 Changed target from LH2 to Carbon. Ask 5uA with raster off for Harp
    scan and Moller measurement.
    Eugene came to do the moller measurement.
    20:28 Moller measurement is finished.
    Joe started the ep measurement
    Jack went down to the Hall to adjust the gas flow.

    00:00 Joe is still doing the ep measurement. FPP straw efficiency is slowly getting
    better.


    _____________________________________________________________________
    Codarun# Harp#
    2548 945

    Time Run Comment
    20:03 2549 VDC and FPP chamber check
    2550 test run
    2551 test run
    2552 test run
    2553 test runs for VDC and FPP chesk.
    _____________________________________________________________________